OpenStudy (is3535):

What is the median of the data set? {10, 15, 14, 14, 10, 10, 8, 18, 11, 12, 17, 16}

OpenStudy (is3535):

@Vocaloid

OpenStudy (is3535):

@Thesmarterone

OpenStudy (black_ninja123):

Median is the middle number

OpenStudy (anonymous):

^that will help you ALOT^

OpenStudy (black_ninja123):

the answer is 13

OpenStudy (is3535):

it 9

OpenStudy (anonymous):

(12+14)/2 = 13

OpenStudy (is3535):

no it 13

OpenStudy (black_ninja123):

You have to put the number from lease to greatest first

OpenStudy (anonymous):

8,10,10,10,11,12,14,14,15,16,17,18

OpenStudy (black_ninja123):

14+12=26 and 26/2 is 13 so @TheOrion101 is right

OpenStudy (anonymous):

Total Numbers: 12 Mean (Average): 12.91667 Median: (12+14)/2 = 13 Mode: 10 Range: 10

OpenStudy (anonymous):

Thanks @Black_ninja123 . Once again, the website I linked above is really useful.

OpenStudy (is3535):

so 13

OpenStudy (black_ninja123):

Np

OpenStudy (black_ninja123):

Yes

OpenStudy (anonymous):

As well as http://www.calculatorsoup.com/calculators/

OpenStudy (is3535):

The ages of people on the camp bus are listed. 10, 15, 14, 8, 18, 11, 12, 12, 10, 10, 17, 16 What is the mode of the ages? A. 10 B. 17 C. 8 D. 12

OpenStudy (black_ninja123):

Also they need to know how to do this without a calc too

OpenStudy (is3535):

i think it 12

OpenStudy (black_ninja123):

Mode is the most repeated

OpenStudy (black_ninja123):

its not 12

OpenStudy (is3535):

10

OpenStudy (black_ninja123):

Yes

OpenStudy (anonymous):

which number is repeated most often?

OpenStudy (anonymous):

Good job.

OpenStudy (black_ninja123):

If you need anymore help just tag us